# Community detection in the hypergraph stochastic block model and reconstruction on hypertrees

### v1

@@ 1.2 Reconstruction on hypertrees
Finally, Theorem 5 is proved by combining both parts. We first take $\epsilon>0$ in the robust reconstruction part such that Theorem 6 holds. Then we take $d$ large enough such that large degree asymptotics works and Eq. (13) holds.
We remark that our proof of Theorem 5 also gives the asymptotic reconstruction threshold for BOHT with $r\geq 7$.
@@
Using Theorem 7, we can prove Theorem 2(ii) by computing a lower bound on $w^{*}(r)$. The proof is deferred to Section 2.5.
@@ 1.3 Organization
In Section 2 we prove our results regarding the BOHT model, including Theorem 5 (tightness of KS for $r\leq 6$), Theorem 6 (robust reconstruction below and at KS), Theorem 7 (asymptotic reconstruction threshold), and their implications for HSBM weak recovery, Theorem 1 (tightness of KS for $r\leq 6$) and Theorem 2(ii) (asymptotic weak recovery lower bound). In Section A we provide missing proofs of lemmas used in Section 2.
In Section B we prove Theorem 4 (non-asymptotic weak recovery upper bound) and Theorem 2(i) (asymptotic weak recovery upper bound).
In Section C we make discussions on related works and further directions.
@@ 2.1 Preliminaries
A binary memoryless symmetric (BMS) channel is a channel $P:\{\pm\}\to\mathcal{Y}$ together with a measurable involution $\sigma:\mathcal{Y}\to\mathcal{Y}$ such that $P(E|+)=P(\sigma(E)|-)$ for all measurable $E\subseteq\mathcal{Y}$. Examples of BMS channels include binary erasure channels (BECs) and binary symmetric channels (BSCs). Furthermore the broadcast channel $B_{r,\lambda}$ (Eq. (8)) and the channels $M_{k}$ (defined as $\sigma_{\rho}\mapsto(T_{k},\sigma_{L_{k}})$) are all BMS channels (with the involution being the global sign flip).
Every BMS channel is equivalent to a mixture of BSC channels. That is every BMS channel $P$ is equivalent to a channel $X\mapsto(\Delta,Z)$ where $\Delta\in\left[0,\frac{1}{2}\right]$ is independent of $X$ and $P_{Z|\Delta,X}=\operatorname{\mathrm{BSC}}_{\Delta}(\cdot|X)$. We use $\Delta_{P}$ to denote the $\Delta$-component of $P$. It is a random variable taking values in the interval $\left[0,\frac{1}{2}\right]$. In this way, BMS channels (up to channel equivalence) are in one-to-one correspondence with distributions on $\left[0,\frac{1}{2}\right]$.
@@
Let $P:\mathcal{X}\to\mathcal{Y}$ and $Q:\mathcal{X}\to\mathcal{Z}$ be two channels with the same input alphabet. We define the $\star$-convolution channel $P\star Q:\mathcal{X}\to\mathcal{Y}\times\mathcal{Z}$ as $(P\star Q)(E\times F|x)=P(E|x)Q(F|y)$ for measurable subsets $E\subseteq\mathcal{Y}$, $F\subseteq\mathcal{Z}$. For $n\in\mathbb{Z}_{\geq 0}$, we use $P^{\star n}$ to denote the $n$-th $\star$-convolution power.
@@
Let $\mathcal{A}$ be a collection of information channels with the same input alphabet $\mathcal{X}$ and $\mu$ be a distribution on $\mathcal{A}$. Then we write $\mathbb{E}_{P\sim\mu}P$ to be the channel which maps $x\in\mathcal{X}$ to $(P,Y)$, where $P\sim\mu$ and $Y\sim P(\cdot|x)$. This channel is called a mixture of $\mathcal{A}$.
@@ 2.2 Large degree asymptotics
Fix $2\leq r\leq 6$. Then $g_{r,\frac{1}{r-1}}(x)<x$ for all $0<x\leq 1$.
Proof of Lemma 10 is deferred to Section A.2.
Proposition 11.
Fix $2\leq r\leq 6$. For any $\epsilon>0$, there exists $d_{0}=d_{0}(r,\epsilon)$ such that for any $d\geq d_{0}$ and $\lambda\in\left[-\frac{1}{2^{r-1}-1},1\right]$ with $(r-1)d\lambda^{2}\leq 1$, we have
$\lim_{k\to\infty}C_{\chi^{2}}(M_{k})<\epsilon.$ (32)
Proof.
By Lemma 10, for any $\epsilon>0$ there exists $\delta>0$ such that $g_{r,\frac{1}{r-1}}(x)\leq x-\delta$ for all $x\in[\epsilon,1]$. Take $d_{0}=d_{0}(r,\delta/2)$ in Prop. 8. Then for any BMS channel $P$ with $C_{\chi^{2}}(P)\geq\epsilon$, we have
@@ 2.3 Robust reconstruction
If $r=2$, then the result follows from the reconstruction threshold for the BOT model ([BRZ95, EKPS00]). In the following assume $r\geq 3$.
@@
Take $\gamma$ large enough such that Lemma 13 and Lemma 14 hold. Take $\epsilon$ such that Eq. (51) has no fixed point for all $x\leq\epsilon$. Such $\epsilon$ exists because the coefficient of $x^{2}$ is lower bounded by $0.4$. Then for any $\gamma$-normal BMS channel $P$ with $C_{\chi^{2}}(P)\leq\epsilon$, robust reconstruction with respect to $P$ is impossible.
@@
2.4 Proof of BOHT results
@@
Take $\epsilon=\epsilon(r)$ in Theorem 6. Take $d_{0}=d_{0}(r,\epsilon)$ in Prop. 11. Fix $d\geq d_{0}$, $\lambda\in\left[-\frac{1}{2^{r-1}-1},1\right]$ such that $(r-1)d\lambda^{2}\leq 1$. By Prop. 11, there exists $k_{0}$ such that $C_{\chi^{2}}(M_{k_{0}})\leq\epsilon$. By Theorem 6, we have
@@
Upper bound. Let $w=w^{*}+\epsilon$. By Lemma 16, it suffices to prove for $(r-1)d\lambda^{2}=w$. By Lemma 9(i) and definition of $w^{*}$, there exists $x_{0}\in(0,1]$ such that $g_{r,w}(x_{0})>x_{0}$. Let $\epsilon=g_{r,w}(x_{0})-x_{0}$. Take $d_{0}=d_{0}(r,\epsilon)$ in Prop. 8. By Prop. 8, for $d\geq d_{0}$ and $(r-1)d\lambda^{2}=w$, for any BMS channel $P$ with $C_{\chi^{2}}(P)\geq x_{0}$, we have
Lower bound. Let $w=w^{*}-\epsilon$. By Lemma 9(i) and definition of $w^{*}$, we have $g_{r,w}(x)<x$ for all $x\in(0,1]$. By the same proof as Prop. 11, for any $\epsilon^{\prime}>0$ there exists $d_{0}=d_{0}(r,\epsilon^{\prime})$ such that $\lim_{k\to\infty}C_{\chi^{2}}(M_{k})<\epsilon^{\prime}$. The rest of the proof is the same as the proof of Theorem 5. ∎
2.5 Proof of HSBM results
Theorem 15 (HSBM-BOHT reduction, [Gu23, Theorem 5.15]).
Consider the model $\operatorname{\mathrm{HSBM}}(n,r,a,b)$ where $r\geq 2$ and $a,b\in\mathbb{R}_{\geq 0}$. Let $d,\lambda$ be as defined in Eq. (1). If reconstruction is impossible for $\operatorname{\mathrm{BOHT}}(r,\lambda,\operatorname{\mathrm{Pois}}(d))$, then weak recovery is impossible for $\operatorname{\mathrm{HSBM}}(n,r,a,b)$.
Proof of Theorem 1.
By Theorem 15 and Theorem 5. ∎
Proof of Theorem 2(ii).
is a polynomial with non-negative coefficients. Therefore for $0\leq x\leq 1$, $\frac{s_{r,w}(x)}{x}$ is maximized at $x=1$, taking value $w2^{r-2}=1$. So $w^{*}(r)\geq w$. Applying Theorem 7 and Theorem 15 we finish the proof. ∎
@@ Acknowledgments
Y.G. is supported by the National Science Foundation under Grant No. DMS-1926686.
@@ A.1 More preliminaries
Let $P:\mathcal{X}\to\mathcal{Y}$ and $Q:\mathcal{X}\to\mathcal{Z}$ be two channels. We say $P$ is more degraded than $Q$, denoted $P\leq_{\deg}Q$, if there exists channel $R:\mathcal{Z}\to\mathcal{Y}$ such that $P=R\circ Q$. If $P$ and $Q$ are BMSs, then it is possible to take $R$ such that it respects the involution structure of BMSs.
For two BMS channels $P$ and $Q$, $P\leq_{\deg}Q$ if and only if there exists a couple between $\Delta_{P}$ and $\Delta_{Q}$ such that $\mathbb{E}[\Delta_{Q}|\Delta_{P}]\leq\Delta_{P}$ almost surely.
@@
Let $R$ be the following channel: on input $y^{n}\in\mathcal{Y}^{n}$, it outputs $y^{n}$ with probability $c$, and outputs a sample from $(P\circ\operatorname{\mathrm{Unif}}(\{\pm\}))^{\times(r-1)}$ with probability $1-c$. It is easy to check that Eq. (62) holds. ∎
Let $P:\{\pm\}\to\mathcal{Y}$ be a BMS channel with discrete target space $\mathcal{Y}$. For $x\in\{\pm\}$, we use $P_{x}$ to denote the distribution $P(\cdot|x)$. The $\theta$-component $\theta_{P}$ is directly related with the distribution $P_{+}$. For $y\in\mathcal{Y}$, define
@@ A.2 Large degree asymptotics
$=\mathbb{E}_{Z\sim\mathcal{N}(0,1)}\frac{1}{2}\left(\tanh(s+\sqrt{s}Z)+\tanh(s-\sqrt{s}Z)\right)$ (67)
$\leq\tanh(s)\leq s.$ (68)
@@
By Lemma 9(i), it suffices to prove the result for $r=6$. Recall function $g$ defined in Eq. (65). Our goal is to show that
@@
Therefore it remains to prove that $g\left(x+2x^{3}+\frac{x^{5}}{5}\right)\leq x-0.006$ for all $x\in S$. We show that this can be done by a rigorous numerical integration. Suppose we would like to evaluate $g(s)$ where $s=x+2x^{3}+\frac{x^{5}}{5}$. Let $n$ be a positive integer to determined later.
@@ A.3 A priori estimates
$x^{\prime}_{k+1}$ $=C_{\chi^{2}}(M_{P,k}^{\times(r-1)}\circ B_{r,\lambda})$ (93)
$=\mathbb{E}_{\theta_{1},\ldots,\theta_{r-1}}C_{\chi^{2}}((B_{\theta_{1}}\times\cdots\times B_{\theta_{r-1}})\circ B_{r,\lambda}).$
@@
Here $D$ denotes either the point distrbution at $d$, or $\operatorname{\mathrm{Pois}}(d)$.
@@
Proof of Lemma 12, Eq. (45) and Eq. (46).
Eq. (45) and Eq. (46) directly follow from Eq. (41), Eq. (42), Eq. (43), Eq. (44). ∎
Thus we have finished the proof of Lemma 12.
@@
Take $k_{0}=\Theta(1/\delta)$ such that $\frac{1}{2}\leq(1-2\delta)^{k_{0}}$, $(1+2\delta)^{k_{0}}\leq 2$. Then for all $0\leq k\leq k_{0}$, we have

### v2

@@ 1.2 Reconstruction on hypertrees
Finally, Theorem 5 is proved by combining Prop. 6 and Theorem 7. We first take $\epsilon>0$ in the robust reconstruction part such that Theorem 7 holds. Then we take $d$ large enough such that large degree asymptotics works and Eq. (14) holds.
Our proof of Theorem 5 also determines the asymptotic reconstruction threshold for BOHT with $r\geq 7$.
@@ 1.3 Organization
In Section 2 we present the proof structure for Theorem 5 (tightness of the KS threshold for $r\leq 6$ and large $d$), which includes the two main parts Prop. 6 (large degree asymptotics) and Theorem 7 (robust reconstruction below and at KS). In Section 3 we prove Theorem 2 (non-asymptotic HSBM weak recovery upper bound).
In Section A we provide missing proofs of lemmas used in Section 2. In Section B we provide proofs of our remaining main results, including Corollary 3 (non-tightness of the KS threshold), Theorem 4 (asymptotic HSBM weak recovery bounds), and Theorem 8 (asymptotic BOHT reconstruction threshold). In Section C we make discussions on related works and further directions.
@@ 2.1 Preliminaries
A binary memoryless symmetric (BMS) channel is a channel $P:\{\pm\}\to\mathcal{Y}$ together with a measurable involution $\sigma:\mathcal{Y}\to\mathcal{Y}$ such that $P(E|+)=P(\sigma(E)|-)$ for all measurable $E\subseteq\mathcal{Y}$. Examples of BMS channels include binary erasure channels (BECs) and binary symmetric channels (BSCs). The broadcast channel $B_{r,\lambda}$ (Eq. (9)) and the channels $M_{k}$ (defined as $\sigma_{\rho}\mapsto(T_{k},\sigma_{L_{k}})$) are all BMS channels (with the involution being the global sign flip).
Every BMS channel is equivalent to a mixture of BSC channels. That is, every BMS channel $P$ is equivalent to a channel $X\mapsto(\Delta,Z)$ where $\Delta\in\left[0,\frac{1}{2}\right]$ is independent of $X$ and $P_{Z|\Delta,X}=\operatorname{\mathrm{BSC}}_{\Delta}(\cdot|X)$. We use $\Delta_{P}$ to denote the $\Delta$-component of $P$. It is a random variable taking values in the interval $\left[0,\frac{1}{2}\right]$. In this way, BMS channels (up to channel equivalence) are in one-to-one correspondence with distributions on $\left[0,\frac{1}{2}\right]$.
@@
Let $P:\mathcal{X}\to\mathcal{Y}$ and $Q:\mathcal{X}\to\mathcal{Z}$ be two channels with the same input alphabet. We define the $\star$-convolution channel $P\star Q:\mathcal{X}\to\mathcal{Y}\times\mathcal{Z}$ as $(P\star Q)(E\times F|x)=P(E|x)Q(F|x)$ for measurable subsets $E\subseteq\mathcal{Y}$, $F\subseteq\mathcal{Z}$. For $n\in\mathbb{Z}_{\geq 0}$, we use $P^{\star n}$ to denote the $n$-th $\star$-convolution power.
@@
Let $\mathcal{A}$ be a collection of information channels with the same input alphabet $\mathcal{X}$ and $\mu$ be a distribution on $\mathcal{A}$. We define $\mathbb{E}_{P\sim\mu}P$ to be the channel that maps $x\in\mathcal{X}$ to $(P,Y)$, where $P\sim\mu$ and $Y\sim P(\cdot|x)$. This channel is called a mixture of $\mathcal{A}$.
@@ 2.2 Large degree asymptotics
Fix $r\leq 6$. Then $g_{r,\frac{1}{r-1}}(x)<x$ for all $0<x\leq 1$.
Proofs of Lemma 10 and Lemma 11 are deferred to Section A.2.
Proof of Prop. 6.
By Lemma 11, for any $\epsilon>0$ there exists $\delta>0$ such that $g_{r,\frac{1}{r-1}}(x)\leq x-\delta$ for all $x\in[\epsilon,1]$. Take $d_{0}=d_{0}(r,\delta/2)$ in Prop. 9. Then for any BMS channel $P$ with $C_{\chi^{2}}(P)\geq\epsilon$, we have
@@ 2.3 Robust reconstruction
Take $\gamma$ large enough such that Lemma 13 and Lemma 14 hold. Take $\epsilon$ such that Eq. (51) has no fixed point for all $x\leq\epsilon$. Such $\epsilon$ exists because the coefficient of $x^{2}$ is upper bounded by $-0.4$. Then for any $\gamma$-normal BMS channel $P$ with $C_{\chi^{2}}(P)\leq\epsilon$, robust reconstruction with respect to $P$ is impossible.
@@
2.4 Finishing the proof
We are now ready to prove Theorem 5.
@@ 2.4 Finishing the proof
Take $\epsilon=\epsilon(r)$ in Theorem 7. Take $d_{0}=d_{0}(r,\epsilon)$ in Prop. 6. Fix $d\geq d_{0}$, $\lambda\in\left[-\frac{1}{2^{r-1}-1},1\right]$ such that $(r-1)d\lambda^{2}\leq 1$. By Prop. 6, there exists $k_{0}$ such that $C_{\chi^{2}}(M_{k_{0}})\leq\epsilon$. By Theorem 7, we have
@@
3 HSBM weak recovery upper bound
In this section we prove Theorem 2. We say an event happens a.a.s. (asymptotically almost surely) if it happens with probability $1-o(1)$ as $n\to\infty$.
@@ 3 HSBM weak recovery upper bound
We first describe our algorithm. Suppose we are given a HSBM instance $G=(V,E)$. We say a partition $Y:V\to\{\pm\}$ is almost balanced if |#$\left|\#Y^{-1}(+)-\frac{n}{2}\right|\leq n^{0.6}$. We say a hyperedge $e\in E$ is in-community with respect to $Y$ if $Y_{v}$s are equal for all $v\in e$. We say a partition $Y:V\to\{\pm\}$ is good if the number of in-community hyperedges is in $\left[\frac{an}{2^{r-1}r}-n^{0.6},\frac{an}{2^{r-1}r}+n^{0.6}\right]$, and bad otherwise. The algorithm looks for a good almost balanced partition. If such a partition exists, return any of them; otherwise report failure.
We say a partition $Y$ is correlated with $X$ (the true partition) if $d_{H}(Y,X)\leq\left(\frac{1}{2}-\epsilon\right)n$, and uncorrelated with $X$ otherwise. We prove that for some $\epsilon>0$, a.a.s. the following are true.
  1. (a)
    There exists a good almost balanced partition.
  2. (b)
    All almost balanced partitions that are uncorrelated with $X$ are bad.
If (a) holds, our algorithm returns a good partition. If in addition (b) holds, the returned partition is correlated with the true partition $X$, and the algorithm succeeds. It remains to prove (a) and (b).
Proof of (a)
We show that $X$ is a.a.s. a good almost balanced partition. By the central limit theorem, $X$ is a.a.s. almost balanced. For such $X$, the number of in-community sets $S\in\binom{V}{r}$ is $\binom{|X|}{r}+\binom{n-|X|}{r}=(1\pm O(n^{-1}))\frac{n^{r}}{2^{r-1}\cdot r!}$. Each such set is an hyperedge independently with probability $\frac{a}{\binom{n}{r-1}}=(1\pm O(n^{-1}))\frac{a(r-1)!}{n^{r-1}}$. By Chebyshev’s inequality, a.a.s. the number of in-community hyperedges is in $\left[\frac{an}{2^{r-1}r}-n^{0.6},\frac{an}{2^{r-1}r}+n^{0.6}\right]$.
Proof of (b)
  1. 1.
    $X$ is almost balanced.
  2. 2.
    The total number of hyperedges $m$ is in $\left[\frac{dn}{r}-n^{0.6},\frac{dn}{r}-n^{0.6}\right]$.
We show that (b) is true conditioned on $m$. We consider the following random hypergraph model $G(n,m,X)$. For each of the $m$ hyperedge, generate a sequence $(x_{1},\ldots,x_{r})$ of $r$ bits according to the following distribution
$p_{x_{1},\ldots,x_{r}}=\left\{\begin{array}[]{ll}\frac{1}{2}\cdot\frac{a}{a-b+2^{r-1}b},&\text{if}~{}x_{1}=\cdots=x_{r},\\ \frac{1}{2}\cdot\frac{b}{a-b+2^{r-1}b},&\text{o.w.}\end{array}\right.$ (58)
Then choose $r$ vertices $v_{1},\ldots,v_{r}$ independently according to $v_{i}\sim\operatorname{\mathrm{Unif}}(X^{-1}(x_{i}))$, and add hyperedge $e=(v_{1},\ldots,v_{r})$ to the hypergraph.
$c_{++}=\left(\frac{1}{4}+\frac{\delta}{2}\right)n\pm O(n^{0.6}),$ $c_{+-}=\left(\frac{1}{4}-\frac{\delta}{2}\right)n\pm O(n^{0.6}),$ (59)
$c_{-+}=\left(\frac{1}{4}-\frac{\delta}{2}\right)n\pm O(n^{0.6}),$ $c_{--}=\left(\frac{1}{4}+\frac{\delta}{2}\right)n\pm O(n^{0.6}).$
Consider the process of generating one hyperedge under $G(n,m,X)$. Let $y_{i}=Y_{v_{i}}$. Then we have
$\mathbb{P}[y_{i}=+|x_{i}=+]=\frac{1}{2}+\delta\pm O(n^{-0.4}),\qquad\mathbb{P}[y_{i}=+|x_{i}=-]=\frac{1}{2}-\delta\pm O(n^{-0.4}).$ (60)
Therefore, under $G(n,m,X)$, each hyperedge is in-community with respect to $Y$ with probability
$~{}\sum_{x_{1},\ldots,x_{r}\in\{\pm\}}p_{x_{1},\ldots,x_{r}}\left(\prod_{i\in[r]}\mathbb{P}[y_{i}=+|x_{i}]+\prod_{i\in[r]}\mathbb{P}[y_{i}=-|x_{i}]\right)$ (61)
By Sanov’s theorem, the probability that there are $\left[\frac{an}{2^{r-1}r}-n^{0.6},\frac{an}{2^{r-1}r}+n^{0.6}\right]$ in-community hyperedges with respect to $Y$ is
By Eq. (5), for $\epsilon>0$ close enough to 00, the last expression is $\exp\left(-n(\log 2+\epsilon^{\prime})\pm o(n)\right)$ for some $\epsilon^{\prime}>0$. Therefore every uncorrelated balanced partition $Y$ is good with probability $\exp\left(-n(\log 2+\epsilon^{\prime})\pm o(n)\right)$. On the other hand, the total number of partitions is $2^{n}$. By union bound, a.a.s. there are no good uncorrelated balanced partitions. This finishes the proof of (b). ∎
@@ Acknowledgments
Y.G. is supported by the National Science Foundation under Grant No. DMS-1926686. We thank anonymous reviewers for helpful comments and suggestions.
@@ A.1 More preliminaries
Let $P:\mathcal{X}\to\mathcal{Y}$ and $Q:\mathcal{X}\to\mathcal{Z}$ be two channels. We say $P$ is more degraded than $Q$, denoted $P\leq_{\deg}Q$, if there exists a channel $R:\mathcal{Z}\to\mathcal{Y}$ such that $P=R\circ Q$. If $P$ and $Q$ are BMSs, then it is possible to take $R$ such that it respects the involution structure of BMSs.
For two BMS channels $P$ and $Q$, $P\leq_{\deg}Q$ if and only if there exists a coupling between $\Delta_{P}$ and $\Delta_{Q}$ such that $\mathbb{E}[\Delta_{Q}|\Delta_{P}]\leq\Delta_{P}$ almost surely.
@@
Let $R$ be the following channel: on input $y^{r-1}\in\mathcal{Y}^{r-1}$, it outputs $y^{r-1}$ with probability $c$, and outputs a sample from $(P\circ\operatorname{\mathrm{Unif}}(\{\pm\}))^{\times(r-1)}$ with probability $1-c$. It is easy to check that Eq. (67) holds. ∎
Let $P:\{\pm\}\to\mathcal{Y}$ be a BMS channel with a discrete target space $\mathcal{Y}$. For $x\in\{\pm\}$, we use $P_{x}$ to denote the distribution $P(\cdot|x)$. The $\theta$-component $\theta_{P}$ is directly related with the distribution $P_{+}$. For $y\in\mathcal{Y}$, define
@@ A.2 Large degree asymptotics
$=\mathbb{E}_{Z\sim\mathcal{N}(0,1)}\frac{1}{2}\left(\tanh(s+\sqrt{s}Z)+\tanh(s-\sqrt{s}Z)\right)$
$\leq\tanh(s)\leq s.$
@@
Recall function $g$ defined in Eq. (70). Because $g$ is strictly increasing on $[0,1]$ and $s_{r,\frac{1}{r-1}}(x)$ is non-decreasing in $r\in\{2,3,4,5,6\}$ for $0\leq x\leq 1$, $g_{r,\frac{1}{r-1}}(x)=g\left(s_{r,\frac{1}{r-1}}(x)\right)$ is non-decreasing in the same range. Therefore it suffices to prove the statement for $r=6$.
Our goal is to show that
@@
Therefore it remains to prove that $g\left(x+2x^{3}+\frac{x^{5}}{5}\right)\leq x-0.006$ for all $x\in S$. We do this using a rigorous numerical integration. Suppose we would like to evaluate $g(s)$ where $s=x+2x^{3}+\frac{x^{5}}{5}$. Let $n$ be a positive integer to determined later.
@@ A.3 A priori estimates
$x^{\prime}_{k+1}$ $=C_{\chi^{2}}\left(M_{P,k}^{\times(r-1)}\circ B_{r,\lambda}\right)$ (96)
$=\mathbb{E}_{\theta_{1},\ldots,\theta_{r-1}}C_{\chi^{2}}((B_{\theta_{1}}\times\cdots\times B_{\theta_{r-1}})\circ B_{r,\lambda}),$
where $B_{\theta}$ denotes the channel $\operatorname{\mathrm{BSC}}_{\frac{1-\theta}{2}}$.
@@
Here $D$ denotes either the point distribution at $d$, or $\operatorname{\mathrm{Pois}}(d)$.
@@
Lemma 12, Eq. (45) and Eq. (46) directly follow from Eq. (41), Eq. (42), Eq. (43), Eq. (44). Thus we have finished the proof of Lemma 12.
@@
Take $k_{0}=\Theta_{r}(1/\delta)$ such that $\frac{1}{2}\leq(1-2C\delta)^{k_{0}}$ and $(1+2C\delta)^{k_{0}}\leq 2$. Then for all $0\leq k\leq k_{0}$, we have